OpenStudy (anonymous):

can (6a^2-5a-6) be factored?

OpenStudy (anonymous):

yes it can

OpenStudy (anonymous):

answer is 18/12 and -8/12

OpenStudy (anonymous):

Delta = -5^(2) - 4 (6)(-6) = 5 + 4(6)(6) Yes it can be factored using the quadratic equation

OpenStudy (anonymous):

so simple ax^2+bx+c to solve this or roots of this eqn is (-b+/- root of(b2-4ac))/2a and the answer is 18/12 and -8/12

OpenStudy (anonymous):

what ever the quadratic formula spits out sub into (x-x1)(x-x2)

OpenStudy (anonymous):

x1 and x2 being the two solutions you acquire and you have a factored equation although it will be a little ugly most likely

OpenStudy (anonymous):

(2a - 3) (3a +2) nevermind it isn't dirty

OpenStudy (anonymous):

Oh buddy a way easier way. Did you learn the x method?

OpenStudy (anonymous):

 The top of the x is b from the basic equation ax^2+bx+c The bottom is a times c The left and right need to add to b and multiply to a times c.